OpenStudy (anonymous):

Find the values of x in the following equation: |9x + 9| = 18

OpenStudy (skullpatrol):

Do you want learn how to work with absolute values?

OpenStudy (anonymous):

I just don't get it. >.<

OpenStudy (skullpatrol):

Let me start with the definition: absolute value: The positive number of any pair of opposite nonzero real numbers is the absolute value of each number in the pair.

OpenStudy (skullpatrol):

This definition is stating that the absolute value is a POSITIVE number.

OpenStudy (skullpatrol):

Let me show you with your question how to apply this definition.

OpenStudy (anonymous):

Ah. I see. Thank you very much for helping me. I appreciate it. Continue, please. :3

OpenStudy (skullpatrol):



OpenStudy (anonymous):

See here http://www.wolframalpha.com/input/?i=abs%28x%29

OpenStudy (skullpatrol):

Can you solve the equation $$9x +9 =18$$

OpenStudy (skullpatrol):

Similarly we use the same argument for the opposite of 18 or -18....

OpenStudy (anonymous):

With 9x + 9 = 18, the "x" stands for 3, right? :o

OpenStudy (skullpatrol):

does 9(3) + 9 =?= 18

OpenStudy (skullpatrol):

27 + 9 =36 =/=18

Parth (parthkohli):

Think of it this way: